# Pub/Sub Message Brokers for GenAI

### 22 Dec 2023 v1

@@ Abstract
In today’s digital world, Generative Artificial Intelligence (GenAI) such as Large Language Models (LLMs) is becoming increasingly prevalent, extending its reach across diverse applications. This surge in adoption has sparked a significant increase in demand for data-centric GenAI models, highlighting the necessity for robust data communication infrastructures. Central to this need are message brokers, which serve as essential channels for data transfer within various system components. This survey aims to delve into a comprehensive analysis of traditional and modern message brokers, offering a comparative study of prevalent platforms. Our study considers numerous criteria including, but not limited to, open-source availability, integrated monitoring tools, message prioritization mechanisms, capabilities for parallel processing, reliability, distribution and clustering functionalities, authentication processes, data persistence strategies, fault tolerance, and scalability. Furthermore, we explore the intrinsic constraints that the design and operation of each message broker might impose, recognizing that these limitations are crucial in understanding their real-world applicability. We then leverage these insights to propose a sophisticated message broker framework—one designed with the adaptability and robustness necessary to meet the evolving requisites of GenAI applications. Finally, this study examines the enhancement of message broker mechanisms specifically for GenAI contexts, emphasizing the criticality of developing a versatile message broker framework. Such a framework would be poised for quick adaptation, catering to the dynamic and growing demands of GenAI in the foreseeable future. Through this dual-pronged approach, we intend to contribute a foundational compendium that can guide future innovations and infrastructural advancements in the realm of GenAI data communication.
@@ 1 Introduction
In the burgeoning field of Generative Artificial Intelligence (GenAI), the computing continuum faces unprecedented challenges in efficiently managing data flows and computational resources. GenAI has predominantly targeted consumer applications, offering applications such as the chatGPT, based on a large language model (LLMs), a type of machine learning (ML) model with a deep neural network. However, a surge in machine-to-machine (M2M) use cases beckons a reevaluation of the supporting communication infrastructure. This survey paper delves into the role of publish/subscribe (pub/sub) message broker systems, considering in particular their emerging role for seamless and scalable data exchange in GenAI applications. We scrutinize contemporary message brokers for their adaptability and efficiency in GenAI contexts, outline existing challenges, and chart promising research avenues for future development.
@@
As a result, current computing continuum platforms, spanning the networks and computational resources from user devices to cloud kokkonen2022autonomy ; motlagh2022edge , face new challenges. These platforms provide support AI models, offering interconnect between their data sources and sinks, and optimising their use of resources in the computing continuum. As GenAI models require and generate ever larger amounts of data wang2023overview , all the while consuming computational resources varying from moderate to massive kokkonen2022autonomy , the computing continuum must offer a dynamic and scalable communication and computation substrate to ensure timely data dissemination and efficient use of resources dustdar2022distributed ; computers12100198 .
Approaches based on the publish/subscribe paradigm (in short, pub/sub) are a preferred choice for a variety of real-world AI applications in the computing continuum. Pub/sub approaches decouple data producers from their consumers, allowing applications to develop components independently, and enhancing system robustness and adaptability eugster2003many . Furthermore, pub/sub makes system design more flexible by increasing the independence between system components with a reliable interconnect.
@@
With the advent and growing popularity of GenAI, there is an increased need for robust and adaptive message brokers, equipped to cater to the specific requirements of GenAI applications. Such brokers must be capable of handling the enormous amounts of data required and generated by GenAI models, as well as adeptly monitoring the processes and the data exchanges for high Quality of Service (QoS) loven2023can .
The primary contributions of this survey are summarized as follows:
@@ 3.2.2 No Priority Support
3.3 Summary on Message Brokers
Clustering support ActiveMQCluster enables scaling the message service to accommodate more clients or connections, effectively handling large message volumes and numerous clients. Monitoring OpenMQmonitor tools are crucial for tracking a message broker’s performance and health, allowing for proactive management, early problem detection, and reliable operation.
Pub/sub support ibmmq9 enables separation of publishers from subscribers, increasing system flexibility. Parallel processing support Kafka2 allows the message broker to handle multiple messages simultaneously, improving throughput and efficiency.
Pull and push support Kafka2 allow flexible and timely message delivery. Reliable delivery support Oracle5 ensures messages are not lost during transit, typically through acknowledgments, retries, or temporary storage until successful delivery.
Persistence support Oracle6 safeguards messages against loss during broker restarts or storage message failures. Authentication support Oracle7 is vital for permitting only authorized users and systems to post or subscribe to messages, especially in systems dealing with sensitive data.
4 Message brokers and GenAI
Pub/sub communication patterns from the perspective of GenAI hold significant potential, providing a variety of benefits that could enhance content delivery, personalization, and user engagement through a variety of factors. With the use of these AI models, such as GPT-3 and its successors, it is possible to provide customized content in real-time, tailored to individual subscriber preferences by automatically generating data such as human-like text according to individual subscriber preferences wang2023overview .
For example, language-based GenAI systems can process subscriber queries and feedback efficiently using natural language understanding. Moreover, they can power chatbots and virtual assistants that enable users to communicate both interactively and intelligently with each other openai2023gpt4 . Additionally, such systems can sum up content, translate it into different languages, and moderate it, enhancing quality and accessibility garcia2023we .
As a result, systems can be designed to provide valuable information to subscribers proactively in the form of automated reporting, anomaly detection, and predictive analytics 10198233 . By automating content verification and validation, GenAI contributes to scalability, resource optimization, and load balancing wang2023overview .
Moreover, computing architectures are evolving to address the rising processing and performance demands of GenAI applications. The GenAI agent model (Fig. 3) contains four parts including the environment, providing input for the agent’s perception component, the brain component, with decision-making capabilities and LLMs as core computational engines, as well as the actuation component, capable of acting upon the environment through its functional capability to produce actions stuart2010artificial ; xi2023rise . The agent is initialised from the users in the environment, providing predefined goals to the agent.
The agent uses actual physical actuators, LLM- based tools, or both to execute tasks in the actuation component. These elements allow agents to interact with and respond to their environments. The LLM-based tools include text generation through text-based tools such as ChatGPT chatgpt2023 . Moreover, agents’ workspaces have been expanded with embodied actions to support their integration and interaction with the physical world. LM-Nav shah2023lm analyzes input commands and the environment, aiming to identify the optimal walk based on a topological graph that is constructed internally. EmbodiedGPT mu2023embodiedgpt enables robots to comprehend and perform motion sequences in physical settings through multimodal visual understanding. Using these non-textual output tools extend the functionality of language models and agent scenarios.
4.1 GenAI for message brokers
GenAI can enhance message broker intelligence and efficiency, specifically through predicting routing decisions (by avoiding busy routes) and filtering of messages tailored to specific criteria dhoni2023exploring . Moreover, using GenAI’s predictive capabilities dhoni2023exploring , the message broker can proactively distribute workloads or services across various nodes with minimal access times.
GenAI’s ability to identify error-inducing patterns dhoni2023exploring enables the message broker to implement automatic corrective measures, increasing system reliability and reducing downtime.
Additionally, GenAI can determine optimal times to scale resources so that resources are utilized optimally dhoni2023exploring . Furthermore, GenAI’s ability to learn from ongoing interactions and explainability results in a continuous feedback loop openai2023 that refines system performance.
4.2 GenAI on message brokers
By contrast, asynchronous communication capability ibmmq9 allows message brokers to significantly improve GenAI task processing efficiency by facilitating interactions that do not require real-time communication. This setup also streamlines real-time data stream processing.
A key component of message brokers is their ability to prevent message loss Oracle4 , which is essential for GenAI to prevent misuse of computational resources or loss of critical data. Moreover, message brokers often feature tools that monitor and manage the flow of messages, a key element of maintaining performance and operational integrity of GenAI.
Furthermore, message brokers facilitate service decoupling ibmmq9 , enabling GenAI to manage growing workloads more effectively. The broker’s ability to distribute tasks to different nodes allows GenAI to achieve a balanced load distribution ibmmq9 , with different components running on different nodes. This is particularly important for GenAI deployments in the computing continuum, providing them with access to local environments with limited computational capacity. In such cases, a message broker can act as an intermediary, bridging sensors with the perception module and conveying actions to the actuators and responses to the user.
Moreover, within in the brain component, a message broker can connect the different components, which often require heavy computational capacity. Fostering collaboration between the different LLM models forming up the decision-making component, message brokers can improve agent performance on complex tasks and allow the offloading of tasks in a balanced way among the available resources, thus enabling for efficient use of computing, communication and storage resources.
However, integrating GenAI into the pub/sub form requires careful planning and customization. GenAI may lead to biased and misleading outcomes. Thus, it is crucial to address privacy and ethical issues while aligning with specific system goals 10198233 . In particular, message brokers need advanced methods to handle the massive amounts of data potentially consumed and generated by GenAI applications.
As an example, while 5G or 6G networks can increase connectivity, model compression methods and model training acceleration may reduce GenAI computational requirements and delays. Furthermore, the efficient resource management process enables workload distribution based on the capacity of each node, thus improving the performance of the distributed system. Moreover, by turning GenAI input and output data into priority-based smart data, timely processing can be further improved. Furthermore, choosing the right models and updating them as data changes improves results. Finally, a continuous monitoring system can detect anomalies and data loss immediately, allowing for corrections and adaption in real time.
In the following sections, we will discuss the advanced methods designed to enhance message brokers’ functionality within the GenAI context, overcoming deployment challenges. By tackling these issues, we aim for seamless integration and optimal performance of GenAI applications.
4.3 Semantic communication
Effective management and real-time analysis of vast data volumes are crucial for developing GenAI applications. This process must be flexible, accommodating various data types with utmost precision. To alleviate strain on communication networks, it’s essential to embed efficient (e.g., semantic) communication mechanism xia2023generative within message brokers. Such integration can enable intelligent, automated feature selection aligned with subscriber needs, streamlining the transmission of data-rich content donta2023governance . Consequently, this approach facilitates smart data transfer, optimizing resource allocation and minimizing latency.
Message brokers equipped with dynamic prioritization capabilities can intelligently identify and route high-priority messages by incorporating semantic communication technology. Prioritization allows message brokers to handle messages based on content along with criteria within message header. This feature not only reduces network congestion but also preserves network resources. Additionally, integrating semantic communication with a broker requires high scalability, processing power, and memory to handle big data, while maintaining robust security features to process sensitive data.
In delay-sensitive applications like healthcare, this integration is vital to assigning priorities based on its deep understanding of data patterns and sensitivities and its subscribers’ specific needs. For example, in the healthcare scenario, GenAI can analyze large amounts of medical data to identify urgent cases, alerting healthcare professionals of critical patient needs or alarming health trends oniani2023military . Similarly, GenAI can prioritize alerts based on the severity and location of incidents at fire stations, ensuring a prompt response.
Furthermore, Apache RocketMQ, with its support for message broadcasting and filtering ApacheRocketMQdocs , has the potential to offer advanced message routing and filtering capabilities by leveraging semantic communication. These capabilities involve interpreting messages’ semantic content, allowing for more precise and context-sensitive handling.
Finally, Anypoint MQ’s intelligent routing AnypointMQ1 combined with semantic filtering allows messages to be identified according to their inherent importance and meaning. It also ensures faster processing, routing of urgent and high priority content, and timely responses.
4.4 Dynamic data and model management
Because GenAI models such as LLMs often process sensitive or personal data, robust security measures are critical. For example, Google Cloud Pub/Sub GoogleCloudPub/Sub incorporates multiple integrated security measures to protect confidential data being transmitted. These measures include authentication and encryption using keys managed by Google, along with the latest Data Encryption Key (DEK) technology, among others GoogleCloudPub/Sub6 .
Moreover, Amazon Kinesis Kinesis offers reliable data transmission methods, utilizing synchronous data replication, checkpointing mechanisms, and error handling strategies Kinesis2 ; Kinesis3 .
4.5 Training acceleration
4.6 Dynamic Model Compression
4.7 Dynamic Orchestration
For instance, Celery can handle, maintain, and schedule distributed tasks across various nodes, thereby enhancing the efficiency of GenAI agents Celerydev ; Celeryintro . Apache ActiveMQ provides efficient resource allocation ApacheActiveMQ . Thus, optimizing Apache ActiveMQ is critical to providing intelligent resource allocation to ensure balanced workloads across different nodes.
Amazon SQS provides a reliable queue service for handling messages that facilitates microservices and distributed systems decoupling SQS1 . Further, the robust architecture of Apache Kafka, which uses clusters of brokers to handle data distribution and partition topics for scalability, makes it excellent for supporting distributed event streaming Kafka1 . Additionally, Azure Service Bus provides advanced scheduling features and message orchestration in distributed environments AzureServiceBus5 ; AzureServiceBus4 .
4.8 Dynamic System Monitoring
MLOps, merging DevOps principles with machine learning (ML), is central to the advancement of ML and AI, streamlining the lifecycle of models from development to deployment. A critical feature within this field is the monitoring of deployed models, crucial for the uninterrupted and reliable operation of message broker systems. This practice enables real-time insights into model metrics, resource usage, and system irregularities, creating a proactive environment for identifying and addressing issues promptly. Furthermore, MLOps facilitates the setting up of automated alerts and triggers, enhancing responsiveness to anomalies and minimizing downtime loven2023can ; kreuzberger2023machine .
A message broker can be embedded with monitoring tools like MLOps and CDM, but it is a complex process and resource hungry. It requires distributed architectures and parallel processing to handle high-intensive tasks effectively and timely. There is a chance of starvation between regular data transmissions and monitoring. However, this integration demands the message broker’s ability to allocate and manage computational resources intelligently. This requires striking an optimal balance between monitoring and regular tasks to gain performance optimization, lower cost, and energy efficiency.
As KubeMQ supports multistage pipelines KubeMQ2 , it could be essential for managing data pipelines within MLOps frameworks. Moreover, HiveMQ and Amazon Kinesis can contribute to MLOps integration by facilitating real-time monitoring and alert systems integration, since they support real-time device monitoring HiveMQ ; Kinesis .
Finally, tracking GenAI model performance can be effectively achieved through integration with IBM MQ, which has robust monitoring and tracing capabilities ibmmq2 . Solace PubSub could potentially assist in identifying and resolving issues related to message routing, delivery, and processing Solace1 , thus facilitating control and modeling activities in MLOps environments.
4.9 Impact of 5G and 6G on message brokers
The advent of 5G technology has significantly influenced message brokers by enabling faster data rates, reduced latency, and enhanced connectivity and reliability. 5G efficiently manages vast message volumes, crucial for real-time applications, and supports numerous devices simultaneously without triggering network congestion 8767268 .
The anticipated 6G evolution promises further enhancements, such as higher throughput, even lower latency, and AI integration at the network level loven2019edgeai ; peltonen20206g ; peltonen2022many . These advancements will empower message brokers to dynamically optimize message routing, adapt to fluctuating network conditions, assure consistent reliability, and process extraordinary data volumes. This capacity is essential for more sophisticated applications necessitating stringent data security. To that end, 6G is projected to fortify security protocols, a critical aspect of protecting future message broker systems porambage2019sec .
5 Conclusion
In this paper, we summarized recent message brokers by highlighting their features and limitations. We categorized these message brokers using criteria such as their open-source availability, integrated monitoring tools, message prioritization mechanisms, parallel processing capabilities, reliability, clustering and distribution capabilities, authentication processes, data persistence strategies, fault tolerance, and scalability.
Further, we explored the limitations of the message brokers, and reflected on requirements for a robust message broker framework that would meet GenAI specifications. To that end, we discussed possible approaches for enhancing the effectiveness of message brokers to support the rapid development of GenAI applications. These approaches include, for example, model training acceleration, model compression, resources orchestration, large-scale data processing, and monitoring systems performance. Furthermore, we discussed the impact of 5G and 6G technologies on message brokers.
By examining the features, challenges, and potential approaches for future research on message brokers, our study will provide the basis for developing a flexible communication system that can distribute data efficiently, provide high-quality service, manage resources intelligently, and meet the needs of ever-growing GenAI applications.
@@ 5 Conclusion
Message delivery guarantees are limited ActiveMQDocument .
@@
Scaling is challenging ActiveMQPolicies .
@@
Works on small and medium-scale applications ActiveMQPolicies .
@@
Compatibility issues between versions ApacheQpidReleases .
@@
Detects failures and assigns messages to different brokers ApacheQpidConn .
Low latency ApacheQpid3 .
Supports multiple authentication schemes ApacheQpid1 .
@@
from malicious activity ApacheQpidbroker .
@@
intensive and strain brokers RabbitMQcmemory .
@@
Clustering has few features and is complicated RabbitMQcheck .
Message size is limited to 512MB donta2022survey ; RabbitMQcons .
Runs on all major operating systems RabbitMQplatforms .
Has good documentation RabbitMQdocs .
Works with C, C++, .NET, and Python RabbitMQtools .
@@
Supports multiple messaging protocols RabbitMQprotocol .
Offers several built-in exchange types RabbitMQfeature .
@@
messages flooding RabbitMQflow .
@@
can be complicated Celeryintro .
Overall complexity Celeryintro .
Monitoring and management are challenging Celerymonitor .
Number of connections is limited by 10 connections Celerylimit .
@@
,such as starting, stopping, and restarting worker processes Celerydev .
Functions as a task queue Celeryintro .
Focuses on real-time processing Celerydev .
Supports task scheduling Celeryintro .
Supports multi-message brokers Celeryintro .
Integrates with multi-web frameworks Celeryintro .
@@
Supports AMQP, MQTT, STOMP message protocols JBosspro .
Supports transactions JBosstra .
@@
and access control JBosscon .
Easy integration with other JBoss and Java $E$ components JBoss31 .
@@
Setup is complex OpenMQoverview .
@@
Loosely-coupled architecture OpenMQoverview .
@@
Manual configuration Gearmandoc .
Monitoring tools are limited Gearmanprotocol .
Used by LiveJournal, Yahoo!, and Digg GearmanIntro .
Multi-languages support Gearmandoc .
@@
Buffer size is limited to max 64KB Enduro/X1 ; Mavimax2022Enduro/X1 .
Cluster nodes number is limited to max 32 nodes Mavimax2022Enduro/X .
@@
are limited to max 32 Mavimax2022Enduro/X .
@@
executed within the callback Enduro/X1 .
Distributed transaction processing Mavimax2022Enduro/X .
@@
Resource intensive for maintenance HiveMQ3 .
@@
Suitable for mission-critical applications HiveMQ1 .
@@
Supports MQTT bridging EMQX2 .
Supports data integration EMQX1 .
@@
Complex configuration and deployment Pulsar2 .
@@
An index-based storage system Pulsar1 .
@@
Resource intensive Kafka1 .
Provides a data backlog Kafka1 .
Complex Kafka1 .
Supports topic (log) compaction & distributed event streaming Kafka1 .
@@
Supports multiple data formats Kafka1 .
@@
consumer groups Kafka1 .
Supports replication and partitioning of data Kafka1 .
Supports deployment in different environments Kafka1 .
@@
Message size is limited to max 4MB ApacheRocketMQ2 .
Message sending retries is limited to max 3 times ApacheRocketMQ2 .
Supports message broadcasting, tracking, filtering, and retrying ApacheRocketMQdocs .
Low latency ApacheRocketMQ1 .
Maintains the order of messages ApacheRocketMQdocs .
Supports multi-protocols ApacheRocketMQoverview .
Supports multiple programming languages ApacheRocketMQoverview .
@@
Limited security EclipseMosquitto1 .
@@
QoS support EclipseMosquitto4 .
Topic-based message filtering EclipseMosquitto5 .
Supports logging and debugging EclipseMosquitto4 .
Supports functioning as a bridge EclipseMosquitto4 .
Supports dynamic restart configuration EclipseMosquitto4 .
@@
Brokerless messaging platform zeromq1 .
@@
Carries messages across IPC, TCP, TPIC, and multicast zeromq1 .
@@
separated from a cluster ApacheNiFidoc .
@@
deleted ApacheNiFidoc .
Complex configuration ApacheNiFiConfigure .
Provides a data flow framework ApacheNiFidoc .
@@
data size ApacheNiFidoc .
@@
production of more data than a queue can handle ApacheNiFidoc .
Supports buffering of all queued data ApacheNiFioverview .
Integrates and processes multiple data sources ApacheNiFioverview .
@@
Caps the number of channels per connection AblyRealtime2 .
@@
240 AblyRealtime2 .
Message size limited to 16KB AblyRealtime2 .
Number of queues limited to 5 AblyRealtime2 .
Queue length limited to 10,000 AblyRealtime2 .
@@
Supports streaming data AblyRealtime1 .
@@
Supports at-least once data processing SamZa1 .
Real-time data processing with low latency SamZa1 .
Easy to integrate SamZa1 .
@@
Master-less clustered messaging protocol VerneMQ1 .
@@
Supports transactions and recovery is built-in NServiceBus2 .
@@
Low clustering capabilities kestrel1 .
@@
Number of items in the queue is limited to 500 kestrel1 .
@@
Data loss with server crash NSQ2 .
Messages are unordered NSQ2 .
Limited persistence NSQ2 .
No message recovery NSQ2 .
Lacks replication NSQ2 .
@@
Efficient handling of high-volume& real-time data streams NSQ ; NSQ1 .
NATS NATS Both Message size is limited to max 64MB NATS2 .
@@
Offers persistence with ”at-least-once” and ”exactly-once” NATS2 .
@@
and tasks queue use cases KubeMQ2 .
@@
messages ibmmq2 .
Controls undelivered messages ibmmq4 .
Multi-APIs support ibmmq5 .
Allows applications to be decoupled ibmmq3 .
Easy to deploy on various platforms ibmmq3 .
@@
High scale-up cost SQS2 .
@@
Message ordering is not guaranteed SQS4 .
@@
and 14 days SQS4 .
Cloud-based web service SQS1 .
@@
systems, and serverless applications SQS1 .
@@
Messages are delivered at least once SQS4 .
@@
May experience resource failure MSMQ6 .
Limitation on message size MSMQ6 .
@@
Multi-protocols support MSMQ1 .
Tracks and deletes expired messages MSMQ5 .
Manages distributed brokers MSMQ4 .
Supports remote access MSMQ4 .
Effective routing MSMQ2 .
Provides guaranteed message delivery MSMQ2 .
Provides a store and forward mechanism MSMQ4 .
Supports transactions MSMQ3 .
@@
Message Queue Oracle1
@@
Resource intensive as the number of messages increases Oracle3 .
@@
increases Oracle3 .
Size of message is limited to max 70MB Oracle1 .
Supports transactions Oracle1 .
@@
Excludes fault tolerance of the server TIBCO2 .

### 21 Feb 2024 v2

# Towards Message Brokers for Generative AI: Survey, Challenges, and Opportunities

@@ Abstract
In today’s digital world, Generative Artificial Intelligence (GenAI) such as Large Language Models (LLMs) is becoming increasingly prevalent, extending its reach across diverse applications. This surge in adoption has sparked a significant increase in demand for data-centric GenAI models, highlighting the necessity for robust data communication infrastructures. Central to this need are message brokers, which serve as essential channels for data transfer within various system components. This survey aims to delve into a comprehensive analysis of traditional and modern message brokers, offering a comparative study of prevalent platforms. Our study considers numerous criteria including, but not limited to, open-source availability, integrated monitoring tools, message prioritization mechanisms, capabilities for parallel processing, reliability, distribution and clustering functionalities, authentication processes, data persistence strategies, fault tolerance, and scalability. Furthermore, we explore the intrinsic constraints that the design and operation of each message broker might impose, recognizing that these limitations are crucial in understanding their real-world applicability. Finally, this study examines the enhancement of message broker mechanisms specifically for GenAI contexts, emphasizing the criticality of developing a versatile message broker framework. Such a framework would be poised for quick adaptation, catering to the dynamic and growing demands of GenAI in the foreseeable future. Through this dual-pronged approach, we intend to contribute a foundational compendium that can guide future innovations and infrastructural advancements in the realm of GenAI data communication.
@@ 1 Introduction
In the burgeoning field of Generative Artificial Intelligence (GenAI), the computing continuum faces unprecedented challenges in efficiently managing data flows and computational resources. GenAI has predominantly targeted consumer applications, offering applications such as the ChatGPT chatgpt , a conversational AI agent based on a large language model (LLMs), a type of machine learning (ML) model with a deep neural network. However, a surge in machine-to-machine (M2M) use cases, coupled with the increasing possibility of relying more and more on decentralized, distributed, and edge-based Large Language Models (LLMs), beckons a reevaluation of the supporting communication infrastructure. This reevaluation is necessitated by the evolving demands for higher bandwidth, lower latency, and more robust data processing capabilities that these advanced applications require shen2024large ; tarkoma2023ai . This survey paper delves into the role of publish/subscribe (pub/sub) message broker systems, considering in particular their emerging role for seamless and scalable data exchange in GenAI applications. We scrutinize contemporary message brokers for their adaptability and efficiency in GenAI contexts, outline existing challenges, and chart promising research avenues for future development.
@@
As a result, current computing continuum platforms, spanning the networks and computational resources from user devices to cloud kokkonen2022autonomy ; motlagh2022edge , face new challenges. These platforms provide support AI models, offering interconnect between their data sources and sinks, and optimising their use of resources in the computing continuum. As GenAI models require and generate ever larger amounts of data wang2023overview , all the while consuming computational resources varying from moderate to massive kokkonen2022autonomy , the computing continuum must offer a dynamic and scalable communication and computation substrate to ensure timely data dissemination and efficient use of resources dustdar2022distributed ; Praveen2023exploring .
Pub/sub approach is equally useful alongside other paradigms across the computing continuum for a wide range of AI applications, including smart cities 8782370 , healthcare 7098706 , and many other domains. This approach decouples data producers from their consumers, allowing applications to develop components independently, and enhancing system robustness and adaptability eugster2003many . Furthermore, pub/sub makes system design more flexible by increasing the independence between system components with a reliable interconnect.
@@
As GenAI continues to evolve, parallels can be drawn with the historical trajectory of IoT systems, where the emergence of robust and adaptive message brokers marked a significant evolutionary step. These message brokers became a de facto paradigm, primarily because they addressed critical challenges associated with scalability, real-time data processing, and the integration of heterogeneous devices and platforms redondi2019towards . Similarly, in the context of GenAI, the anticipation of an analogous development is not without merit. The complexity and volume of data that GenAI applications demand, coupled with the necessity for high-quality service monitoring and data exchange processes, suggest that a transition towards more sophisticated message brokering solutions may be inevitable. Such solutions would not only have to manage the large data throughput but also ensure adaptability, reliability, and efficiency in dynamic GenAI ecosystems. Reflecting on the IoT evolution, the motivations for this shift include the need to support scalable communication, facilitate interoperability among diverse systems, and uphold stringent Quality of Service (QoS) standards, which are likely to be paralleled in the GenAI domain  loven2023can .
However, it is critical to acknowledge that the evolution towards more sophisticated message brokering solutions, specifically tailored to accommodate GenAI application needs, does not come without its challenges.
Bearing all this in mind, the primary contributions of this survey are summarized as follows:
@@ 3.2.2 No Priority Support
Message delivery guarantees are limited ApacheActiveMQ .
@@
Scaling is challenging ApacheActiveMQ .
@@
Works on small and medium-scale applications ApacheActiveMQ .
@@
Compatibility issues between versions ApacheQpid .
@@
Detects failures and assigns messages to different brokers ApacheQpid .
Low latency ApacheQpid .
Supports multiple authentication schemes ApacheQpid .
@@
from malicious activity ApacheQpid .
@@
intensive and strain brokers RabbitMQ .
@@
Clustering has few features and is complicated RabbitMQ .
Message size is limited to 512MB donta2022survey ; RabbitMQ .
Runs on all major operating systems RabbitMQ .
Has good documentation RabbitMQ .
Works with C, C++, .NET, and Python RabbitMQ .
@@
Supports multiple messaging protocols RabbitMQ .
Offers several built-in exchange types RabbitMQ .
@@
messages flooding RabbitMQ .
@@
can be complicated Celery .
Overall complexity Celery .
Monitoring and management are challenging Celery .
Number of connections is limited by 10 connections Celery .
@@
,such as starting, stopping, and restarting worker processes Celery .
Functions as a task queue Celery .
Focuses on real-time processing Celery .
Supports task scheduling Celery .
Supports multi-message brokers Celery .
Integrates with multi-web frameworks Celery .
@@
Supports AMQP, MQTT, STOMP message protocols JBossCons .
Supports transactions JBossCons .
@@
and access control JBossCons .
Easy integration with other JBoss and Java $E$ components JBossCons .
@@
Setup is complex OpenMQ .
@@
Loosely-coupled architecture OpenMQ .
@@
Manual configuration Gearman .
Monitoring tools are limited Gearman .
Used by LiveJournal, Yahoo!, and Digg Gearman .
Multi-languages support Gearman .
@@
Buffer size is limited to max 64KB Mavimax2022Enduro/X1 .
Cluster nodes number is limited to max 32 nodes Mavimax2022Enduro/X1 .
@@
are limited to max 32 Mavimax2022Enduro/X1 .
@@
executed within the callback Mavimax2022Enduro/X1 .
Distributed transaction processing Mavimax2022Enduro/X1 .
@@
Resource intensive for maintenance HiveMQ2 .
@@
Suitable for mission-critical applications HiveMQ .
@@
Supports MQTT bridging EMQX .
Supports data integration EMQX .
@@
Complex configuration and deployment Pulsar .
@@
An index-based storage system Pulsar .
@@
Resource intensive Kafka .
Provides a data backlog Kafka .
Complex Kafka .
Supports topic (log) compaction & distributed event streaming Kafka .
@@
Supports multiple data formats Kafka .
@@
consumer groups Kafka .
Supports replication and partitioning of data Kafka .
Supports deployment in different environments Kafka .
@@
Message size is limited to max 4MB ApacheRocketMQ .
Message sending retries is limited to max 3 times ApacheRocketMQ .
Supports message broadcasting, tracking, filtering, and retrying ApacheRocketMQ .
Low latency ApacheRocketMQ .
Maintains the order of messages ApacheRocketMQ .
Supports multi-protocols ApacheRocketMQ .
Supports multiple programming languages ApacheRocketMQ .
@@
Limited security EclipseMosquitto .
@@
QoS support EclipseMosquitto .
Topic-based message filtering EclipseMosquitto .
Supports logging and debugging EclipseMosquitto .
Supports functioning as a bridge EclipseMosquitto .
Supports dynamic restart configuration EclipseMosquitto .
@@
Brokerless messaging platform zeromq .
@@
Carries messages across IPC, TCP, TPIC, and multicast zeromq .
@@
separated from a cluster ApacheNiFi .
@@
deleted ApacheNiFi .
Complex configuration ApacheNiFi .
Provides a data flow framework ApacheNiFi .
@@
data size ApacheNiFi .
@@
production of more data than a queue can handle ApacheNiFi .
Supports buffering of all queued data ApacheNiFi .
Integrates and processes multiple data sources ApacheNiFi .
@@
Caps the number of channels per connection AblyRealtime .
@@
240 AblyRealtime .
Message size limited to 16KB AblyRealtime .
Number of queues limited to 5 AblyRealtime .
Queue length limited to 10,000 AblyRealtime .
@@
Supports streaming data AblyRealtime .
@@
Supports at-least once data processing SamZa .
Real-time data processing with low latency SamZa .
Easy to integrate SamZa .
@@
Master-less clustered messaging protocol VerneMQ .
@@
Supports transactions and recovery is built-in NServiceBus1 .
@@
Low clustering capabilities kestrel .
@@
Number of items in the queue is limited to 500 kestrel .
@@
Data loss with server crash NSQ .
Messages are unordered NSQ .
Limited persistence NSQ .
No message recovery NSQ .
Lacks replication NSQ .
@@
Efficient handling of high-volume& real-time data streams NSQ .
NATS NATS Both Message size is limited to max 64MB NATS1 .
@@
Offers persistence with ”at-least-once” and ”exactly-once” NATS1 .
@@
and tasks queue use cases KubeMQ .
@@
messages ibmmq9 .
Controls undelivered messages ibmmq9 .
Multi-APIs support ibmmq9 .
Allows applications to be decoupled ibmmq9 .
Easy to deploy on various platforms ibmmq9 .
@@
High scale-up cost SQS .
@@
Message ordering is not guaranteed SQS .
@@
and 14 days SQS .
Cloud-based web service SQS .
@@
systems, and serverless applications SQS .
@@
Messages are delivered at least once SQS .
@@
May experience resource failure MSMQ .
Limitation on message size MSMQ .
@@
Multi-protocols support MSMQ .
Tracks and deletes expired messages MSMQ .
Manages distributed brokers MSMQ .
Supports remote access MSMQ .
Effective routing MSMQ .
Provides guaranteed message delivery MSMQ .
Provides a store and forward mechanism MSMQ .
Supports transactions MSMQ .
@@
Message Queue Oracle
@@
Resource intensive as the number of messages increases Oracle .
@@
increases Oracle .
Size of message is limited to max 70MB Oracle .
Supports transactions Oracle .
@@
Excludes fault tolerance of the server TIBCO1 .